AB INITIO CALCULATION OF UO2

\underline{QIANG CHANG}, and RUSSELL M. PITZER, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Ab initio spin-orbit configuration interaction calculations are performed on the UO_{2} molecule, which has been observed in matrix isolation and vacuum ultraviolet spectroscopic studies. UO_{2} is isoelectronic with P uO_{2}^{2+}, which has a 5fδ^{1}5fφ^{1} ^{3}H_{4g} ground state. We find, as have other workers, that UO_{2} has a 5fφ^{1}7sσ^{1} ^{3}Φ_{2} ground state with a moderate mixture of 5fδ^{1}7sσ^{1} ^{3}Δ_{2}. the UO distance is found to be 1.786 Ȧ. The ^{3}H_{4g} state is only slightly higher in energy. Oscillator strengths have been calculated for electronic transitions.